Solve by substitution method

y = x + 4
3y – 5x = 6

Answer by rfer(16322)

3(x+4)-5x=6
3x+12-5x=6
-2x=-6
x=3
------------------
3y-5*3=6
3y-15=6
3y=21
y=7
-------------
7=3+4